OpenStudy (anonymous):

Suppose S and T are mutually exclusive events. Find P(S or T) if P(S) = 6/11 and P(T) = 1/10. (a) 3/55 (b) 71/110 (c) 1/3 (d) 71/21

OpenStudy (anonymous):

what do mean by this plz explain your needs

OpenStudy (ciarán95):

If two events, say S and T, are mutually exclusive/independent of each other, the probability of S occuring is exactly the same no matter if T occurred or not, and vice versa (e.g. the outcome from tossing a coin and the outcome from rolling a die). In such cases, we can say that: \[P(S and T) = P(S)P(T)\] In the question, you're asking for the probability of S or T occurring. To be honest, I'm not sure how you would go about working that one out, if you could at all, especially given the fact that you are told so little information. If you wanted to work out P(S and T), given that P(S) = 6/11 and P(T) = 1/10, then you would multiply these two values together to give you the answer: \[P(SandT) = (\frac{ 6 }{ 11 })(\frac{ 1 }{ 10 }) = ?\] I'm assuming here that there was a typo in the question and that 'or' should have been 'and'. I've worked it out and the answer you get is one of the options given above, which you can also do yourself. Hope that helps you out, but apologies if it doesn't and the question was correct. :)